# Grounding Emotion Recognition with Visual Prototypes: VEGA - Revisiting CLIP in MERC

### v1

@@ 1. Introduction
Importantly, while most prior approaches that leverage CLIP focus on its textual modality to align features, typically mapping audio or visual inputs to text-derived embeddings (e.g., prompts such as “this is a sad face”), we deliberately reverse this trend. These prior methods use text prompts because CLIP’s textual encoder offers a powerful, zero-shot compatible representation space and is easier to manipulate through language. However, textual descriptions of emotions are often ambiguous, abstract, and culturally variable, whereas emotional expressions in human faces are more concrete, universal, and visually grounded (Fehr and Russell, 1984; Shaver et al., 1987; Hu et al., 2025, 2024a; Wei et al., 2024; Kollias et al., 2024c). By anchoring fusion around CLIP’s visual embeddings, we embrace a more perceptually grounded and psychologically aligned representation of emotion, closer to how humans perceive affective states.
@@ 2. Related Work
Multimodal Emotion Recognition in Conversations (MERC). While traditional ERC has focused on textual input (Poria et al., 2019; Pereira et al., 2022), its limited emotional expressiveness has driven the development of MERC, which integrates text, audio, and visual modalities (Li et al., 2021; Zhang and Li, 2023b; Kollias et al., 2024f, g, 2025b; Zadeh et al., 2018; Hu et al., 2021b; Yoon et al., 2018). Early fusion approaches (Li et al., 2021) simply concatenate features, but often struggle with modality imbalance and noise. More advanced strategies leverage attention (Yoon et al., 2018; Zhao et al., 2023), memory (Zadeh et al., 2018), context-aware modeling (Majumder et al., 2019; Hu et al., 2021b), and semantic refinement (Zhang et al., 2023; Zhang and Li, 2023b) to better capture inter-modal and temporal dependencies. However, most existing models remain data-driven, lacking structured semantic priors or emotion-level prototypes to guide multimodal alignment (Hu et al., 2025; Kollias et al., 2024b, 2017a, 2020a, 2020c, g; Kollias, 2023; Kollias et al., 2024e; Psaroudakis and Kollias, 2022; Kollias et al., 2021b, 2019a, 2019b; Kollias and Zafeiriou, 2019; Kollias, 2022; Kollias et al., 2023e, f, d, 2024f, 2020b; Zafeiriou et al., 2017; Hu et al., 2024a; Kollias et al., [n. d.]; Kollias and Zafeiriou, 2021a, b; Kollias et al., 2023a, c; Arsenos et al., 2023; Salpea et al., 2022; Arsenos et al., 2024b; Karampinis et al., 2024; Arsenos et al., [n. d.]; Miah et al., 2024; Arsenos et al., 2024a; Kollias et al., 2018b; Gerogiannis et al., 2024; Kollias et al., 2024d, a; Kollias and Zafeiriou, 2020a, 2018d, 2018b; Kollias et al., 2018a, 2017b; Tagaris et al., 2017, 2018; Kollias et al., 2016, 2015, 2025b; Arsenos et al., 2022; Kollias et al., 2023b; Kollias and Zafeiriou, 2020b, 2018c, 2018a; Chamalke Senadeera et al., 2024; Kollias et al., 2021a, 2025a; Senadeera et al., 2024; Karampinis et al., 2025).

### v2

@@ 1. Introduction
Importantly, while most prior approaches that leverage CLIP focus on its textual modality to align features, typically mapping audio or visual inputs to text-derived embeddings (e.g., prompts such as “this is a sad face”), we deliberately reverse this trend. These prior methods use text prompts because CLIP’s textual encoder offers a powerful, zero-shot compatible representation space and is easier to manipulate through language. However, textual descriptions of emotions are often ambiguous, abstract, and culturally variable, whereas emotional expressions in human faces are more concrete, universal, and visually grounded (Fehr and Russell, 1984; Shaver et al., 1987; Hu et al., 2025, 2024a; Wei et al., 2024; Kollias et al., 2024a). By anchoring fusion around CLIP’s visual embeddings, we embrace a more perceptually grounded and psychologically aligned representation of emotion, closer to how humans perceive affective states.
@@ 2. Related Work
Multimodal Emotion Recognition in Conversations (MERC). While traditional ERC has focused on textual input (Poria et al., 2019; Pereira et al., 2022), its limited emotional expressiveness has driven the development of MERC, which integrates text, audio, and visual modalities (Li et al., 2021; Zhang and Li, 2023b; Kollias et al., 2024b, c, 2025; Zadeh et al., 2018; Hu et al., 2021b; Yoon et al., 2018). Early fusion approaches (Li et al., 2021) simply concatenate features, but often struggle with modality imbalance and noise. More advanced strategies leverage attention (Yoon et al., 2018; Zhao et al., 2023), memory (Zadeh et al., 2018), context-aware modeling (Majumder et al., 2019; Hu et al., 2021b), and semantic refinement (Zhang et al., 2023; Zhang and Li, 2023b) to better capture inter-modal and temporal dependencies. However, most existing models remain data-driven, lacking structured semantic priors or emotion-level prototypes to guide multimodal alignment.